OpenStudy (anonymous):

Use the rational expression below to answer the following questions a) what values of x are restrictions on the variable b) simplify the rational expression

OpenStudy (anonymous):

here's the rational expression and problems, its for a finals study guide

OpenStudy (joannablackwelder):

Values of x that make any denominator 0 (any time throughout the simplification) are restrictions.

OpenStudy (joannablackwelder):

Any thoughts on what those might be?

OpenStudy (anonymous):

no idea, i took most of this class a year ago, and algebra 1 my freshman year so these concepts are super vague for me

OpenStudy (anonymous):

if i get what youre saying right, its any number in the numberator that can make the denominator 0? i dont remember how to simplify the expressions like this exactly

OpenStudy (johnweldon1993):

It's just any number in general that you can plug in for 'x' that make the denominator = 0 \[\large \frac{x^2 + 3x - 4}{5 - x} \div \frac{x^2 + 5x + 4}{x^2 - 2x - 15}\] What numbers can you plug in here for 'x' would make the denominator of either fraction = 0?

OpenStudy (anonymous):

im not sure, would i just plug in the values i have until i get a denominator or 0 with one of the x values? again its been over a year since ive touched this concept so i'm not sure how to do it

OpenStudy (johnweldon1993):

Well I mean yeah that is 1 way to do it But lets see The denominator of the first fraction is \(\large 5 - x\) so what can you make 'x' that will make that = 0? The denominator of the second fraction is \(\large x^2 - 2x - 15\) so what can 'x' be that would make this = 0?

OpenStudy (anonymous):

to make the denominator on the left size 0 i guess x=5 so i can check that off as a restriction? 5-5=0

OpenStudy (johnweldon1993):

Right...x = 5 is 1 restriction

OpenStudy (anonymous):

so on the other side i have to find x^2-2x-15 and whatever x is has to make all of the come out to 0?

OpenStudy (johnweldon1993):

Right Just factor and make it easier we need 2 numbers that multiply to make -15 and at the same time add to make -2 -5 and 3 come to mind so the factors of that would be \[\large (x - 5)(x + 3)\] And so what can you make 'x' be so that comes out to 0?

OpenStudy (anonymous):

so (5-5)(-3-3) would equal (0)(0)

OpenStudy (anonymous):

or should x be the same in both the factors?

OpenStudy (johnweldon1993):

Nope it can definitely be 2 different numbers (5 - 5)(anything) = 0 so if x = 5 this = 0 (anything)(-3 - 3) = 0 so if x = -3 this = 0 So we have 1 new restriction to cross off...-3

OpenStudy (anonymous):

okay, so those would be the restrictions, 5 and -3 because both of those when plugged in make the denominator 0

OpenStudy (johnweldon1993):

Sorry that was a typo...it should have been (-3 + 3)=0 But still correct

OpenStudy (johnweldon1993):

And yes, that statement is correct But that is just so far...as another poster has stated above, restrictions can appear at ANY point during simplification :) so we need to begin simplifying

OpenStudy (anonymous):

yeah, that's what i really need help with, I dont remember how to simplify these

OpenStudy (johnweldon1993):

Alright Well first...do you know how I got *above* \[\large x^2 - 2x - 15 = (x - 5)(x + 3)\]?

OpenStudy (anonymous):

i know you factored them 5 and three are both factors of 15 ? that's about all i know on that though

OpenStudy (johnweldon1993):

Okay, well no that's pretty much it :) just wanted to make sure! So we have \[\large \frac{x^2 + 3x - 4}{5 -x} \div \frac{x^2 + 5x + 4}{x^2 - 2x - 15}\] I would begin by turning this into multiplication....so to do that...all we do is flip the second fraction and turn this into multiplication.. So \[\large \frac{x^2 + 3x - 4}{5 -x} \times \frac{x^2 -2x - 15}{x^2 + 5x + 4}\] Okay?

OpenStudy (anonymous):

alrighty, i see what you did there

OpenStudy (johnweldon1993):

Good, so now...that WAS considered a simplification...so we need to check if that introduced any more restrictions! The first fraction is still the same But now...the second fraction has a denominator of \(\large x^2 + 5x + 4\) So what can 'x' be that makes that = 0?

OpenStudy (anonymous):

ok so i have to factor that part first right then find the value of x

OpenStudy (johnweldon1993):

So I would factor that as well two numbers that multiply to make 4 and add to make 5.....looks like 4 and 1 So \[\large x^2 + 5x + 4 = (x + 4)(x + 1)\] what can 'x' be that makes it = 0?

OpenStudy (johnweldon1993):

Oh, yeah sorry I jumped the gun there >.<

OpenStudy (anonymous):

so the values of x are -4 and -1 to make that 0 looks like we did the same thing at the same time

OpenStudy (johnweldon1993):

Lol well yes both of those are now our new restrictions :)

OpenStudy (anonymous):

okay, so the multiplication part, thats conisdered simplified, and the 5, -1, -4, -3 are the restrictions ?

OpenStudy (johnweldon1993):

That is considered the FIRST part of the simplification...we still have more to do :D And yes all of those are the correct restrictions!

OpenStudy (anonymous):

whoops last part was an accident, whats the next part of the simplification?

OpenStudy (johnweldon1993):

Now, to simplify this down lets catch up quick We had \[\large \frac{x^2 + 3x - 4}{5 - x} \times \frac{x^2 - 2x - 15}{x^2 + 5x + 4}\] And we have already done some factoring so lets write those out here too \[\large x^2 + 3x - 4 = (x+4)(x-1)\] \[\large x^2 - 2x - 15 = (x+5)(x-3)\] \[\large x^2 + 5x + 4 = (x+4)(x+1)\] There...all caught up!

OpenStudy (anonymous):

yeah, got all that down on paper!

OpenStudy (johnweldon1993):

So now...since we have all the factors...lets just write those into that equation we have up there \[\large \frac{(x+4)(x-1)}{5-x}\times \frac{(x-5)(x+3)}{(x+4)(x+1)}\]

OpenStudy (johnweldon1993):

So...lets just combine these...since this is all just multiplication \[\large \frac{(x+4)(x-1)(x-5)(x+3)}{(5 - x)(x+4)(x+1)}\]

OpenStudy (johnweldon1993):

And now, we just cancel out what we can...whatever is on top and on bottom will cancel \[\large \frac{\cancel{(x+4)}(x-1)(x-5)(x+3)}{(5 - x)\cancel{(x+4)}(x+1)}\]

OpenStudy (anonymous):

alrighty, i get that

OpenStudy (johnweldon1993):

Now, there is just 1 more thing we need to do Is it true that \(\large (5 - x) = -(x - 5)\) ?

OpenStudy (anonymous):

im not sure, wouldnt the second one mean you have to multiply the inside by -1 ? like -(x-5) would be (-x+5) not entirely sure, i could be reading that wrong

OpenStudy (johnweldon1993):

Right you have to multiply the inside by -1 So we HAD 5 - x And what I'm concluding is that -(x - 5) is the same If we distribute the -1....we have -x + 5 which is the same as 5 - x just written in n another order

OpenStudy (anonymous):

okay

OpenStudy (johnweldon1993):

So with that little....proof? I guess lol we know that we can now write \(\large (5 - x)\) as \(\large -(x - 5)\) So in what we have left of our equation \[\large \frac{(x-1)(x-5)(x+3)}{(5-x)(x+1)}\] We can now write that as \[\large \frac{(x-1)(x-5)(x+3)}{-(x-5)(x+1)}\] So now we have 1 more factor that cancels \[\large \frac{(x-1)\cancel{(x-5)}(x+3)}{-\cancel{(x-5)}(x+1)}\] So finally ALL we have left is \[\large - \frac{(x-1)(x+3)}{(x+1)}\]

OpenStudy (anonymous):

whew! that wasn't too hard to follow thankfully, do you think you could do the complex fraction underneath too with me?

OpenStudy (johnweldon1993):

Sure :)

OpenStudy (anonymous):

thanks! fractions are the main thing i really am not good at

OpenStudy (johnweldon1993):

Alright..so \[\huge \frac{\frac{x}{x-3}}{\frac{1}{x} - \frac{1}{x-3}}\] The first thing we wanna do is get a common denominator on the bottom

OpenStudy (anonymous):

okay, so would a common denominator be 3?

OpenStudy (johnweldon1993):

Not quite... So we have *just the bottom part* \[\large \frac{1}{x} - \frac{1}{x-3}\] To get a common denominator here...we would multiply \(\large \frac{1}{x}\times \frac{(x-3)}{(x-3)}\) and we would also multiply \(\large \frac{1}{x-3} \times \frac{x}{x}\) This would make the denominator in both cases x(x-3) right?

OpenStudy (anonymous):

yeah, im following i think

OpenStudy (johnweldon1993):

Thats just a quick way to find a common denominator...multiply the first fraction by the denominator of the second fraction...and vice-versa But okay...so once we evaluate all that out and combine them over the common denominator ...we have \[\large \frac{(x-3) - (x)}{x(x - 3)}\] right?

OpenStudy (anonymous):

right, i get that!

OpenStudy (johnweldon1993):

Okay good...now that was just the bottom part we focused on...lets bring the top back into this \[\huge \frac{\frac{x}{x-3}}{\frac{(x-3)-x}{x(x-3)}}\] Just a quick thing we can obviously see \[\huge \frac{\frac{x}{x-3}}{\frac{(\cancel{x}-3)\cancel{-x}}{x(x-3)}}\] right? so we really just have \[\huge \frac{\frac{x}{x-3}}{\frac{-3}{x(x-3)}}\]

OpenStudy (anonymous):

arighty, i see what you canceled out there :)

OpenStudy (johnweldon1993):

Great! Okay so now, for this next part...let me just write this as \[\large \frac{x}{x-3} \div \frac{-3}{x(x-3)}\] Just so you can see it like that last question Now, we should turn this into multiplication...how?

OpenStudy (anonymous):

okay so you just seperated it into two fractions like we did in the last question, so to make it multiplication you flip the second fraction right?

OpenStudy (johnweldon1993):

Right and right! :D so we have \[\large \frac{x}{x-3} \times \frac{x(x-3)}{-3}\] So then we just combine them again since its just multiplication! \[\large \frac{x\times x(x-3)}{(x-3) \times -3}\] can you simplify it from there? :)

OpenStudy (anonymous):

the (x-3) on the top and bottom cancel out right?

OpenStudy (johnweldon1993):

Correct :)

OpenStudy (anonymous):

then youd be left with x^2/-3 ?

OpenStudy (johnweldon1993):

Perfect! :) or just \(\large -\frac{x^2}{3}\) but either way yes correct! :)

OpenStudy (anonymous):

awesome! i'm trying to look through the rest of this to make sure there's nothing else on the same topic, i think everything else is sequences, probability, and then trig and functions :(

OpenStudy (anonymous):

Thank you so much! i may ask for some help later on if i get stumped with any of the other questions on those topics!

OpenStudy (johnweldon1993):

Absolutely shoot me a message if you need to...I'm not gonna be on for a little bit after this but if I get on and you still need me I'll be there :)

OpenStudy (anonymous):

thanks!